OpenStudy (anonymous):

A customer is quoted a price of $390 for carpet including installation for one bedroom in her house. She wants to use the same carpet in the living room, which has double the length and width of her bedroom. If the cost per square foot is the same for both rooms, what is the total cost for carpeting the bedroom and the living room, in dollars?

OpenStudy (anonymous):

If the area of the bedroom is l*w, the area of the living room is 2l*2w --> 4*l*w.

OpenStudy (anonymous):

but how do I know what the l and the w are?

OpenStudy (anonymous):

Do you need to find out? You do know that the price for "L*W" is 390usd. You also know that the other rooms area was 4*LW. Do you follow? :)

OpenStudy (anonymous):

This is my worst subject lol, I really don't. :c explain it some other way?

OpenStudy (anonymous):

LW=390 4*LW=? Simply insert the price (value) of LW into above (4*LW), what you'll get then is the price of the living room.

OpenStudy (anonymous):

4*390=1560 390+1560=1950

OpenStudy (anonymous):

You got two rooms. Bedroom and living room. Carpet covers the "Area" of the room. So, to find the answer, you need to find the "relative" area of the living room to the area of the bedroom. Since length and width are "twice", area will be four times larger (2L * 2W = 4LW). Therefore, total cost will be $390*5.

OpenStudy (anonymous):

Omg thats it? I feel dumb. Thank youuuuuuu both. <3 :)

OpenStudy (anonymous):

Well...the dumbness is in the problem! :)